OpenStudy (pokemon23):

My last math question Solve for I in terms of R and E: R=E/T My last math question Solve for I in terms of R and E: R=E/T @Mathematics

OpenStudy (pokemon23):

show all work

OpenStudy (agreene):

there is no "I" to solve for...

OpenStudy (pokemon23):

ops sorry R=E/I

OpenStudy (agreene):

\[R=\frac{E}{I}\] multiply both sides by I \[R*I=E\] divide by R \[I=\frac{E}{R}\]

OpenStudy (unklerhaukus):

I=E/R